OpenStudy (anonymous):

Convert the given exponential function to the form indicated. f(t) = 3.1(1.007)^t; f(t) = Q0e^kt

OpenStudy (anonymous):

f(t) = 3.1* e^(1.007t)

OpenStudy (anonymous):

Nope, it doesn't work.

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Why? How it's possible not working?

OpenStudy (anonymous):

OH , I see!

OpenStudy (anonymous):

f(t) = 3.1* e^(.007t)

OpenStudy (anonymous):

It works now!

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

And, it still doesn't work.

OpenStudy (anonymous):

3.1(1.007)^t initial = 3.1 r = .007 That why P. e^(rt) = 3.1 e^(.007t)

OpenStudy (anonymous):

If it doesn't work, then you need to double check your calculation!

OpenStudy (anonymous):

Aren't you suppose to use this formula f(t) = Q0e^kt